OpenStudy (amy0799):

d/dx m(x) where m(x) is the tangent line to curve x^2+xy-y^2 = 1 at point (2,3)

OpenStudy (amistre64):

the slope of a line is its derivative ...

OpenStudy (amistre64):

d/dx (tangent line to a curve) = y', or slope of the curve

OpenStudy (amistre64):

right?

OpenStudy (amy0799):

yes, so I just need to take the derivative of x^2+xy-y^2 = 1?

OpenStudy (amistre64):

as i see it, yes

OpenStudy (amy0799):

i asked my other friends, and they took the 2nd derivative. so I'm now confused...

OpenStudy (amistre64):

\[y_{tan}=f'(a)(x-a)+f(b)\] \[\frac d{dx}y_{tan}=f'(a)\]

OpenStudy (amistre64):

the derivative of a line, is its slope

OpenStudy (amistre64):



OpenStudy (amistre64):

regardless of the curve ... the function we are taking the derivative of is a line

OpenStudy (amy0799):

\[m'(x)=\frac{ -2x-y }{ x-2y }\] is this correct?

OpenStudy (amistre64):

maybe .. now you gonna make me work lol x^2+xy-y^2 = 1 2xx'+x'y+xy' -2yy' = 0 2 3 2 3 4+3+2y' -6y' = 0 (2-6)y' = -7

OpenStudy (amistre64):

i tend to have less error when i plug in and then solve for y'

OpenStudy (amy0799):

i got 1.75

OpenStudy (amistre64):

7/4 yes

OpenStudy (amy0799):

thank you!

OpenStudy (amistre64):

good luck